### v1

# Physics-Guided Multi-Fidelity DeepONet for Data-Efficient Flow Field Prediction

Abstract
This study presents an enhanced multi-fidelity deep operator network (DeepONet) framework for efficient spatio-temporal flow field prediction, with particular emphasis on practical scenarios where high-fidelity data is scarce. We introduce several key innovations to improve the framework’s efficiency and accuracy. First, we enhance the DeepONet architecture by incorporating a merge network that enables more complex feature interactions between operator and coordinate spaces, achieving a 50.4% reduction in prediction error compared to traditional dot-product operations. We further optimize the architecture through temporal positional encoding and point-based sampling strategies, achieving a 7.57% improvement in prediction accuracy while reducing training time by 96% through efficient sampling and automatic mixed precision training. Building upon this foundation, we develop a transfer learning-based multi-fidelity framework that leverages knowledge from pre-trained low-fidelity models to guide high-fidelity predictions. Our approach freezes the pre-trained branch and trunk networks while making only the merge network trainable during high-fidelity training, preserving valuable low-fidelity representations while efficiently adapting to high-fidelity features. Through systematic investigation, we demonstrate that this fine-tuning strategy not only significantly outperforms linear probing and full-tuning alternatives but also surpasses conventional multi-fidelity frameworks by up to 76%, while achieving up to 43.7% improvement in prediction accuracy compared to single-fidelity training. The core contribution lies in our novel time-derivative guided sampling approach, which strategically selects high-fidelity training points based on temporal dynamics identified by pre-trained low-fidelity models. This physics-guided sampling strategy demonstrates remarkable effectiveness: it maintains prediction accuracy equivalent to models trained with the full dataset while requiring only 60% of the original high-fidelity samples. Through comprehensive experiments across multiple resolutions and dataset sizes, we show that our enhanced framework offers a practical solution for industrial applications where high-fidelity data acquisition is costly and time-consuming.
Keywords Deep operator networks  $\cdot$ Multi-fidelity modeling  $\cdot$ Transfer learning  $\cdot$ Time-derivative guided sampling  $\cdot$ Small dataset  $\cdot$ Spatio-temporal flow prediction
1 Introduction
In the era of scientific machine learning (SciML), accurate and efficient prediction of spatio-temporal flow fields plays a crucial role across various engineering disciplines, particularly in the development of digital twins that require real-time flow field predictions for virtual replicas of physical systems. While traditional numerical approaches like computational fluid dynamics (CFD) offer high accuracy, their computational demands make them impractical for real-time applications, where rapid prediction and flexible adaptation to different flow parameters are essential. This has led to the emergence of various deep learning approaches, including multilayer perceptrons (MLPs) [30, 35], convolutional neural networks (CNNs) [10, 19, 7, 5, 9, 4], and graph neural networks (GNNs) [23, 38, 12, 37], as surrogate models for flow field prediction. However, these architectures face inherent limitations: MLPs lack spatial structure awareness and cannot flexibly handle varying mesh resolutions, requiring retraining for each new mesh configuration—a significant drawback for digital twins that must adapt to different system configurations. CNNs are constrained by fixed grid resolutions, limiting their applicability to irregular domains. GNNs, while capable of handling unstructured meshes, require extensive memory resources to store and process mesh connectivity information, making them computationally intensive for large-scale problems [22].
The deep operator network (DeepONet) [16], introduced to learn mappings between function spaces, offers a compelling alternative for spatio-temporal flow field prediction tasks by overcoming the limitations of traditional methods [8, 1, 17, 26, 25]. Its unique architecture, consisting of separate branch and trunk networks, enables prediction at arbitrary spatial locations without being constrained by specific grid resolutions or mesh connectivities. This resolution-invariant property makes DeepONet particularly suitable for flow field predictions where data may come from simulations with varying mesh configurations. However, the standard DeepONet architecture may not be ideally designed to capture the complex dynamics and intricate features of spatio-temporal behavior [1, 13, 11, 15]. To fully exploit DeepONet’s potential for spatio-temporal flow field prediction, we recognize the need for specialized architectural refinements and training methodologies tailored to fluid flow applications. By systematically developing and integrating enhancements to the core DeepONet framework, we aim to significantly improve both the accuracy and computational efficiency of flow field predictions through: 1) expanding the architecture with additional networks for more flexible feature interactions between operator and coordinate spaces, 2) addressing spectral bias through strategic positional encoding implementation, and 3) implementing efficient point-based sampling techniques with automatic mixed precision training to reduce computational overhead.
However, training DeepONet remains challenging, especially when high-fidelity data is limited due to computational constraints. In this context, multi-fidelity modeling has emerged as a powerful approach in SciML, effectively leveraging abundant low-fidelity data alongside limited high-fidelity data to enhance prediction accuracy while maintaining computational efficiency [6, 20, 36, 39, 34, 33]. Existing multi-fidelity DeepONet frameworks [18, 2, 3] predominantly employ coupled architectures, where low-fidelity (LF) and high-fidelity (HF) networks are trained and utilized together. This coupling approach fundamentally requires identical query points for both branch inputs (function queries) and trunk inputs (spatio-temporal coordinates) across different fidelity levels. This severely restricts DeepONet’s primary advantage: the flexibility to leverage operators at arbitrary points. It forces researchers to design low-fidelity and high-fidelity experiments with identical branch inputs and trunk inputs beforehand—a requirement that is highly inefficient and often impractical in industrial applications where data collection follows operational constraints rather than organized experimental design. This coupled architecture introduces significant computational limitations, requiring sequential two-network inference processes that increase memory requirements and computational costs. The inefficiency scales poorly with multiple fidelity levels, as each additional fidelity layer requires its own network, creating cascading overhead particularly problematic when integrated with physics-informed methods, where automatic differentiation should propagate through all architectures.
Due to this inherent coupling and the availability of matching LF and HF datasets, these approaches naturally adopt residual learning architectures. In these frameworks, LF DeepONet is trained specifically on LF data to produce baseline predictions, while a second separate network, residual DeepONet, is trained to model the discrepancy between the low-fidelity and high-fidelity outputs (i.e., $y_{HF}=y_{LF}+\hat{y}_{residual}$). More fundamentally, residual approaches assume that low-fidelity simulations capture essential physical phenomena with less accuracy—an assumption that can be challenging in fluid dynamics, where low-resolution simulations may struggle to fully represent critical structures like vortex formations or boundary layer effects. When LF and HF data differ in complex ways, residual DeepONet model must focus on modeling these complicated residual patterns, which is highly inefficient compared to directly predicting the HF values. This inefficiency stems from forcing the network to learn fundamentally missing physics rather than leveraging the strengths of end-to-end learning. The residual network must not only correct numerical inaccuracies but also compensate for entirely absent physical phenomena in the baseline LF DeepONet model, essentially requiring it to learn more complex mappings than a direct prediction approach would need.
Beyond these architectural limitations, fundamental methodological challenges exist in effectively leveraging the relationship between LF and HF data during DeepONet training. While Xu et al. [32] proposed a two-phase approach pre-training a LF DeepONet before training a residual DeepONet for high-fidelity predictions—which removes the requirement of identical LF and HF datasets due to its sequential transfer learning approach—their framework still necessitates two separate networks during both training and inference. This decoupling of training between LF and HF DeepONets allows for more flexibility in dataset design, but continues to suffer from computational inefficiency as both models must be employed during inference. Additionally, it lacks a systematic mechanism to identify critical regions where high-fidelity training would be most beneficial, resulting in potentially inefficient use of the pre-trained low-fidelity model and valuable high-fidelity data. Combined with the inherent drawbacks of residual-based architectures discussed earlier, these issues highlight the need for a fundamentally different approach to multi-fidelity operator learning.
A more refined approach to multi-fidelity DeepONet for spatio-temporal flow field prediction would incorporate four key innovations: (1) a unified network structure that eliminates the need for coupled use of low-fidelity and high-fidelity DeepONets during inference, reducing computational overhead and memory requirements; (2) preservation of DeepONet’s inherent query flexibility by removing the requirement for low-fidelity and high-fidelity datasets to share identical function inputs or spatio-temporal query points; (3) a physics-guided sampling strategy that identifies regions of significant flow dynamics using pre-trained low-fidelity models, strategically concentrating high-fidelity data collection in these dynamically critical areas to maximize information gain while minimizing computational cost; and (4) comprehensive evaluation of the relationship between dataset sizes and multi-fidelity model performance, with particular emphasis on strategies for minimizing high-fidelity data requirements while maintaining prediction accuracy.
To address these challenges comprehensively, we present a systematic investigation that directly tackles all four aspects through enhanced DeepONet architectures and efficient training strategies for spatio-temporal flow prediction, with a particular focus on developing practical multi-fidelity frameworks. Our key contributions can be organized into two main thrusts.
Our first step involves designing targeted enhancements to the DeepONet architecture for more accurate spatio-temporal flow field predictions:
  1. 1.
    We propose an expanded DeepONet architecture incorporating a merge network for complex feature interactions between operator (branch net) and coordinate spaces (trunk net). Through extensive experiments, we demonstrate that our element-wise multiplication merge strategy achieves a 50.4% reduction in prediction error compared to traditional dot-product operations, while maintaining similar model parameters. This significant improvement highlights the importance of non-linear feature processing in capturing intricate flow dynamics.
  2. 2.
    We address spectral bias in DeepONet’s coordinate processing through strategic positional encoding implementation. Our comprehensive analysis reveals that selective temporal coordinate encoding achieves a 7.57% improvement in prediction accuracy, while applying encoding to all spatio-temporal coordinates can lead to instability. This investigation provides crucial insights into optimizing coordinate representation while maintaining training stability.
  3. 3.
    We leverage DeepONet’s unique point-based prediction capability to develop efficient training protocols. Our approach demonstrates that training with a subset of spatial points significantly reduces computational overhead while maintaining prediction accuracy—challenging the assumption that more training points always lead to better results. Combined with automatic mixed precision training, we achieve a remarkable 96% reduction in training time (from 8.672 to 0.350 hours) while maintaining or even enhancing prediction accuracy, establishing new benchmarks for practical DeepONet implementation.
Second, we develop and validate a novel multi-fidelity DeepONet (MF-DeepONet) framework that efficiently leverages both low- and high-fidelity data:
  1. 1.
    We introduce a transfer learning-based MF-DeepONet framework that fundamentally departs from sequential, two-network approaches like those in [18, 2, 3]. By strategically transferring knowledge from pre-trained low-fidelity models to a single unified network for high-fidelity predictions, our approach completely eliminates both the computational overhead of maintaining coupled networks and the restrictive constraint of matching query points between fidelity levels. This critical advancement preserves DeepONet’s fundamental capability to evaluate operators at arbitrary points while removing the need for aligned branch and trunk inputs across fidelity levels—enabling seamless integration of heterogeneous data from different sources and resolutions. In direct comparison with conventional multi-fidelity approaches proposed by Lu et al. [18], our framework demonstrates dramatically superior performance with up to 76% improvement in prediction accuracy while significantly reducing computational requirements during both training and inference.
  2. 2.
    Through systematic investigation of different transfer learning configurations—including fine-tuning (where only the merge network is trainable while branch and trunk networks are frozen), full-tuning (where all network components are retrained), and linear probing (where only the final layer is trainable)—we demonstrate that the selective fine-tuning approach significantly outperforms alternatives. Our extensive experimentation with various low-fidelity resolutions (16×16, 32×32, 64×64) and dataset sizes (50-300 samples) reveals that fine-tuning achieves up to 43.7% improvement in prediction accuracy compared to single-fidelity training, while maintaining computational efficiency by preserving valuable representations learned from low-fidelity data.
  3. 3.
    We propose a physics-guided sampling strategy based on temporal derivatives that efficiently identifies critical regions requiring high-fidelity data. This intuitive approach demonstrates remarkable effectiveness: our method reduces prediction error by 10.02-20.73% compared to traditional uniform sampling methods using the same amount of data. Furthermore, using only 60% of high-fidelity samples with the same low-fidelity data, we achieve similar prediction accuracy compared to models trained with the full high-fidelity dataset without sampling. These findings represent a significant breakthrough for industrial applications where high-fidelity data acquisition is prohibitively expensive or time-consuming, providing a principled methodology for dramatically reducing computational requirements without sacrificing prediction accuracy.
The remainder of this paper is organized as follows. Section 2 presents the theoretical foundations of our enhanced DeepONet architecture and multi-fidelity framework, including the merge network design and time-derivative guided sampling strategy. Section 3 describes the flow field data generation process with varying initial conditions and fidelity levels. Section 4 evaluates the architectural enhancements to DeepONet, analyzing the effects of different merge processes, positional encoding, and efficient training techniques. Section 5 investigates our MF-DeepONet framework, comparing different transfer learning approaches and demonstrating performance with limited high-fidelity data. Section 6 explores the time-derivative guided sampling strategy and quantifies its effectiveness in reducing high-fidelity data requirements. Finally, Section 7 summarizes our findings and discusses implications for future research.
2 Methodology
2.1 DeepONet Architecture Enhancement
2.1.1 Enhanced DeepONet with Merge Network
The original DeepONet consists of two main components: a branch network and a trunk network. The branch network maps the input function information (e.g., initial conditions, boundary conditions, or other function characterizations) to a latent representation $\mathbf{B}(u)\in\mathbb{R}^{p}$, where $p$ is the dimension of the latent space. The trunk network takes spatio-temporal coordinates, $(x,y,t)$ in this study, as input and outputs latent space $\mathbf{T}(y)\in\mathbb{R}^{p}$. In the standard DeepONet formulation, these two latent representations are combined through a simple dot product operation to produce the final prediction: $f(u)(y)=\mathbf{B}(u)\cdot\mathbf{T}(y)$.
For complex spatio-temporal flow field predictions, we hypothesize that this simple dot product operation may not provide sufficient non-linear modeling capacity. While Lu et al. [16] demonstrated that incorporating a bias term in the dot product slightly improves the model’s performance in their applications, the effectiveness of this modification in complex flow field prediction remains to be verified. Therefore, we propose enhancing the architecture with a third component—a merge network—and investigate four strategies for merging branch and trunk network outputs, ranging from simple dot products to more complex non-linear operations through proposed merge networks:
  • Merge Type 0 (without bias): Simple dot product from the original DeepONet:
    Final $\text{Final output}=\mathbf{B}(u)\cdot\mathbf{T}(y)$ (1)
  • Merge Type 0 (with bias): Dot product with learnable bias for enhanced flexibility [16]:
    Final $\text{Final output}=\mathbf{B}(u)\cdot\mathbf{T}(y)+b$ (2)
  • Merge Type 1: Element-wise multiplication with additional non-linear processing by merge network:
    Final $\text{Final output}=\text{MergeNet}(\mathbf{B}(u)\odot\mathbf{T}(y))$ (3)
  • Merge Type 2: Concatenation with additional non-linear processing by merge network:
    Final $\text{Final output}=\text{MergeNet}([\mathbf{B}(u);\mathbf{T}(y)])$ (4)
The introduction of MergeNet in Types 1 and 2, as visually represented in Fig. 1, is intended to address a critical limitation of the original DeepONet architecture by enabling more complex non-linear interactions between branch and trunk network outputs. While the simple dot product operation in Type 0 may suffice for simpler operator learning tasks, more sophisticated non-linear processing of branch and trunk outputs can enhance complex spatio-temporal flow field predictions. The merge network provides this crucial non-linear modeling capacity by further processing the combined features through additional neural network layers, enabling the model to capture intricate relationships and dynamic patterns that may be missed by the basic dot product operation alone.
Figure 1: Schematic of two types of merging processes proposed in this study: Type 1 (element-wise) and Type 2 (concatenation).
2.1.2 Enhanced DeepONet with Positional Encoding
Positional encoding (PE) has been widely used in various deep learning architectures, particularly in transformers and attention-based models, to incorporate spatial or temporal information into the input representations [29, 21]. In the context of DeepONet, the trunk net processes coordinate information which can suffer from spectral bias, limiting its ability to capture high-frequency components and intricate spatio-temporal dependencies [31, 28]. By mapping coordinates to a higher-dimensional space beyond the original $(x,y,t)$ representation, PE is expected to enhance the network’s ability to learn complex flow dynamics and temporal dependencies critical for accurate spatio-temporal flow field prediction.
Specifically, when the PE transforms input coordinate $t$ to a vector $\gamma(t)$ using sinusoidal functions,
$\gamma(t)=$ $[a_{1}\cos(2\pi B_{1}^{T}t),a_{1}\sin(2\pi B_{1}^{T}t),\ldots,a_{m/2}\cos(2\pi B_{m/2}^{T}t),a_{m/2}\sin(2\pi B_{m/2}^{T}t)]^{T}$ (5)
where $B_{j}\in\mathbb{R}$ are frequency vectors sampled from a normal distribution and scaled by a factor $\sigma$, $a_{j}$ are amplitude factors (set to 1 in our implementation), and $m$ is the mapping size determining the dimension of the encoded features. This transformation, visualized in Figure 2, maps the input to a higher-dimensional space, enabling better representation of high-frequency functions.
Figure 2: Schematic of the positional encoding process: this example shows the process where only temporal coordinate ($t$) is encoded.
The PE implementation includes two key hyperparameters:
  • Scale Factor ($\sigma$): Controls the frequency range of the encoding, affecting the model’s ability to capture different scales of variation.
  • Mapping Size ($m$): Determines the dimension of the encoded features, controlling the expressiveness of the encoding.
Additionally, we introduce a flexible option that determines the scope of encoding:
  • PE($\mathbf{t}$): PE is applied only to the temporal coordinate in the trunk network input, preserving the original spatial coordinates.
  • PE($\mathbf{x,y,t}$): PE is applied to all coordinates (spatial and temporal) in the trunk network input, providing a uniform encoding scheme.
2.2 Multi-Fidelity DeepONet Framework with Physics-Guided Sampling Strategy
2.2.1 Multi-Fidelity DeepONet Framework
We propose a multi-fidelity DeepONet (MF-DeepONet) framework that efficiently leverages both low-fidelity and high-fidelity data through knowledge transfer. Our two-phase approach first trains a LF DeepONet on abundant low-fidelity data, then strategically transfers these pre-trained networks (branch, trunk, and merge nets) to the high-fidelity model (HF DeepONet). By freezing the branch and trunk parameters (see Figure 3), we preserve the fundamental flow physics captured during low-fidelity training. During the high-fidelity training phase, only the merge network remains trainable, allowing it to learn the specific refinements necessary for accurate high-fidelity predictions. Importantly, the merge network parameters transferred from Phase 1 serve as initial values that are further optimized using the high-fidelity dataset.
Figure 3: Two-phase training process of the MF-DeepONet framework. In Phase 1, all components are trained on low-fidelity data. In Phase 2, all networks are transferred but branch and trunk networks are frozen, while only the merge network is fine-tuned with high-fidelity data.
A key advantage of our approach is its computational efficiency during inference. While training involves two distinct phases, once training is complete, only the final HF DeepONet model in Phase 2 is needed for predictions. This creates a significant computational advantage compared to conventional coupled multi-fidelity frameworks that require both low and high-fidelity networks to be executed in sequence during inference. Our decoupled design effectively eliminates this computational overhead, resulting in substantially faster inference times and reduced memory requirements during deployment—critical factors for real-time applications and resource-constrained environments. Furthermore, this decoupled architecture enables the use of flexible LF and HF datasets, as it doesn’t require corresponding data pairs between fidelity levels.
To comprehensively understand the impact of different knowledge transfer strategies, we will explore three distinct transfer learning approaches within our framework. While our primary proposed MF-DeepONet strategy employs fine-tuning (where only the merge network is trainable while branch and trunk networks are frozen), we additionally investigate [14]: (1) full-tuning, where all network components (branch, trunk, and merge) are trainable, allowing complete adaptation to high-fidelity data; and (2) linear probing, the most restrictive approach that freezes all layers except the final layer of the merge network. By comparing these approaches, we demonstrate the strength of our fine-tuning strategy while providing valuable insights into the trade-offs between preserving low-fidelity knowledge and adapting to high-fidelity features for multi-fidelity modeling of flow fields using DeepONet.
2.2.2 Time-Derivative Sampling Strategy
After training the low-fidelity (LF) DeepONet model, we aim to efficiently utilize the available high-fidelity (HF) dataset for training the HF DeepONet. While DeepONet’s point-based architecture already offers flexibility in training with either complete or subsampled spatial points (note that effects of sub-sampling will be investigated in Section 4.2), we propose further enhancing this capability through physics-guided sampling. Specifically, we introduce a time-derivative sampling technique that strategically identifies dynamically important regions in the flow field. This approach leverages the temporal derivative information from the pre-trained LF model to guide the selection of HF training points, potentially enabling more efficient training of HF DeepONet by focusing computational resources on regions with significant flow dynamics.
The complete procedure for selecting HF training points using time-derivative guided sampling is formally described in Algorithm 1. This physics-guided approach first computes the temporal derivatives at each spatial point from the HF dataset using the pre-trained LF model through automatic differentiation (lines 4-6). For each HF spatial location, we calculate a score based on the average magnitude of temporal derivatives predicted by the LF model (lines 7-9). Points with larger temporal derivatives—indicating regions of significant flow dynamics—receive higher sampling probabilities after score transformation and normalization (lines 10-11). Rather than deterministically selecting the highest-scoring points, we employ probabilistic sampling using probability defined by these score values, which introduces beneficial stochasticity while still favoring dynamically important regions. A fraction of training points (determined by ratio $r$) is selected based on this probabilistic approach (line 13), while the remaining points are sampled uniformly to maintain spatial coverage (line 14). By strategically focusing computational resources on dynamically important regions while preserving randomness, this approach can achieve superior prediction accuracy with fewer HF samples compared to random sampling without considering any physics. Figure 4 visually demonstrates our time-derivative guided sampling strategy, showing how high-fidelity training points are strategically selected based on a combination of derivative-weighted probability distributions and uniform random sampling to maintain spatial coverage.
Algorithm 1 Time-derivative guided sampling strategy
1:Input: HF dataset spatial coordinates $\mathbf{x}_{HF}$, pre-trained LF model, sampling ratio $r$
2:Output: Selected spatial points for HF DeepONet training
3:Note: $\hat{\omega}^{LF}(\mathbf{x}_{HF})$ denotes the vorticity field predicted by the LF model at coordinates $\mathbf{x}_{HF}$
4:for each spatial point $\mathbf{x}_{HF}$ do
5:     Compute $\frac{\partial\hat{\omega}^{LF}(\mathbf{x}_{HF})}{\partial t}$ using automatic differentiation
6:end for
7:for each spatial point $\mathbf{x}_{HF}$ do
8:     $\text{score}(\mathbf{x}_{HF})=\frac{1}{T}\sum_{t=1}^{T}|\frac{\partial\hat{\omega}^{LF}(\mathbf{x}_{HF})}{\partial t}|$
9:end for
10:Transform scores using power-law: $p(\mathbf{x}_{HF})=\text{score}^{2}(\mathbf{x}_{HF})$
11:Normalize: $p(\mathbf{x}_{HF})=\frac{p(\mathbf{x}_{HF})}{\sum p(\mathbf{x}_{HF})}$
12:Among $N$ spatial points in the HF dataset for training HF DeepONet:
13: Sample $N\cdot r$ points using multinomial distribution with probabilities $p(\mathbf{x}_{HF})$
14: Sample remaining points uniformly at random
Figure 4: Time-derivative guided sampling demonstration: background shows probability distribution based on temporal derivatives, with black dots representing points selected using this distribution and white circles showing randomly sampled points for spatial coverage.
3 Data Generation: Flow Fields with Varying Initial Conditions
The training datasets comprise both LF and HF flow field simulations of incompressible Navier-Stokes equations in a two-dimensional periodic domain $[0,1m]\times[0,1m]$. The numerical solution is obtained using a pseudo-spectral solver that employs fast Fourier transforms for spatial discretization, with a 2/3 dealiasing rule to prevent aliasing errors. The solver implements a semi-implicit time integration scheme where advection terms are treated explicitly while diffusion terms are handled implicitly.
First, for the HF dataset, we generate 100 training samples and 50 test samples on a uniform grid of 128×128 resolution. Each sample is initialized with different combinations of two parameters ($var_{1}$, $var_{2}$) that define the initial velocity field components:
$v_{x}(x,y,t=0)$ $=-\sin(var_{1}\pi y)$ (6)
$v_{y}(x,y,t=0)$ $=\sin(var_{2}\pi x)$
where $var_{1}$ and $var_{2}$ are randomly sampled from uniform distributions in the range $\mathcal{U}[1,3]$. The Reynolds number is fixed at $Re=1,000$ for all simulations. The temporal evolution of the flow field is computed with a time step of $\Delta t=2.5\times 10^{-5}[s]$, solving until $t=1s$. While the solver operates at this fine temporal resolution, the flow field data is saved at a coarser interval of $\Delta t_{train}=5\times 10^{-3}[s]$ for training purposes. This output frequency is selected to capture the relevant flow dynamics while maintaining computational efficiency. Importantly, these simulations feature no external forcing after $t=0$, resulting in gradually dissipating flow fields where vorticity magnitude decreases over time. This dissipative behavior creates challenging temporal dynamics that are particularly difficult to predict accurately as the flow evolves further from its initial state. Throughout this study, we focus on predicting the vorticity field, a derived quantity that captures the local rotation in the fluid and presents a more challenging prediction target than primary variables like velocity components. Figure 5 displays the vorticity field at $t=0.5s$ for four distinct initial conditions. The marked diversity in the flow patterns, resulting from variations in $var_{1}$ and $var_{2}$, underscores the complex nature of the dataset.
(a) Example 1
(b) Example 2
(c) Example 3
(d) Example 4
Figure 5: Vorticity fields at $t=0.5s$ for four distinct initial conditions, each defined by a different combination of $var_{1}$ and $var_{2}$. This diversity in the observed patterns highlights the complexity and challenges of predicting the flow evolution, particularly as dissipative effects intensify over time.
For the LF datasets, we generate 300 training samples for each of three different grid resolutions: 16×16, 32×32, and 64×64. As illustrated in Figure 6, the grid resolution significantly impacts the accuracy of flow field predictions. The coarsest resolution (16×16) shows severe limitations in capturing flow evolution: while it approximates the macroscopic flow patterns at $t=0.25s$, the flow field becomes notably distorted from $t=0.5s$ onward compared to higher-resolution simulations. This suggests that data from such coarse resolutions might be inadequate or even detrimental in a multi-fidelity framework. However, from 32×32 resolution upward, the simulations begin to capture the essential macroscopic flow properties visible in the HF (128×128) case. This observation suggests that LF data from these intermediate resolutions (32×32 and 64×64) could effectively support HF predictions in our multi-fidelity framework.
Figure 6: Effects of grid resolution on vorticity field evolution with fixed parameters $var_{1}=3$ and $var_{2}=1$ (contours represent vorticity values).
4 Improvement of DeepONet Architecture for Spatio-Temporal Flow Field Prediction
In this section, we focus on enhancing single-fidelity DeepONet performance using only high-fidelity data (128×128 resolution). We systematically investigate the effects of different merge processes, spatial point sampling approaches, and positional encoding strategies to establish the most effective base architecture. This optimized DeepONet architecture will serve as the foundation for both LF and HF models in our multi-fidelity framework presented in Section 5.
4.1 Effects of Different Merge Process of Branch/Trunk Nets
We first investigate the effectiveness of different strategies for combining outputs from the branch and trunk networks elaborated in Section 2.1.1. To ensure a fair comparison, we carefully adjust the network architectures to maintain similar numbers of learnable parameters across different configurations (approximately 300K parameters). For the basic dot-product variants, Merge Type 0 with and without bias, we employ more complex branch and trunk networks to match the parameter count of models with Type 1 and 2 networks. Across all DeepONet configurations, the branch network consistently takes two input parameters ($var_{1}$ and $var_{2}$) defining initial conditions, while the trunk network processes three spatial-temporal coordinates ($x$, $y$, and $t$). For detailed network architectures, please refer to Table 1. All models are trained for 800 epochs using the Adam optimizer with an initial learning rate of $10^{-3}$, GeLU activation functions, and minibatch size of 1 on an NVIDIA A6000 GPU. To account for training stochasticity and ensure reproducible results, all experiments in this paper are repeated at least three times with different random seeds, and we report the averaged values of these runs.
Table 1: Comparison of different merge strategies for DeepONet
Merge Strategy Network Architecture Parameters Training hour Test Error
Type 0 (Dot-Product w/o bias)
Branch: 2-128×6
Trunk: 3-64-128×6
306,304 6.197 34.447
Type 0 (Dot-Product w/ bias)
Branch: 2-128×6
Trunk: 3-64-128×6
306,305 6.453 34.546
Type 1 (Element-wise)
Branch: 2-128
Trunk: 3-64-128
Merge: 128×6-64-32-16-1
300,289 8.672 17.079
Type 2 (Concatenation)
Branch: 2-128
Trunk: 3-64-128
Merge: 256-128×5-64-32-16-1
292,033 8.399 23.060
The results in Table 1 demonstrate two key findings. First, incorporating merge networks significantly improves prediction accuracy compared to two Type 0 approaches suggested by Lu et al. [16]. While the original dot-product architecture was successful for simpler operator learning tasks in other studies, our results show its limitations in complex flow field predictions: the test error with dot-product operations (both with and without bias) remains above 34, indicating substantial prediction inaccuracies. In contrast, the introduction of merge networks dramatically reduces the test error to 17-23, despite requiring longer training times (approximately 8.5 hours compared to 6.2 hours) due to the additional processing of branch and trunk outputs through the distinct merge network. This stark improvement in accuracy—reducing test error from 34.447 to 17.079, a 50.4% improvement—highlights that additional non-linear processing of combined features is crucial for capturing complex spatio-temporal flow dynamics. The merge network effectively acts as a learnable non-linear mapping that can adapt to intricate relationships between the latent representations from branch and trunk networks, a capability that simple dot-product operations alone cannot provide.
Second, among the enhanced architectures, Type 1 (element-wise multiplication) outperforms Type 2 (concatenation), achieving a test error of 17.079 compared to 23.060. This advantage stems from the structured interaction in Type 1, which more naturally aligns with well-established techniques such as modal decomposition (e.g., proper orthogonal decomposition) and spectral expansions, both of which have demonstrated strong performance in CFD.
  1. 1.
    Proper Orthogonal Decomposition (POD) Perspective: Element-wise multiplication in Type 1 architectures directly mirrors the fundamental structure of POD, where complex flow fields are reconstructed through coefficients multiplied by basis modes [27]. In DeepONet, this correspondence is explicit: the branch network generates coefficients, while the trunk network produces basis functions—creating a natural factorization that resembles POD’s coefficient-times-basis-function formulation. This structural alignment with POD, which has proven highly effective in CFD for reduced-order modeling and flow reconstruction [4, 10], explains why Type 1’s element-wise multiplication outperforms Type 2’s less structured concatenation approach.
  2. 2.
    Spectral Methods Perspective: Type 1 approach also aligns with spectral methods, where solutions are represented as weighted sums of basis functions [15]. Here, the trunk network acts as a global basis generator, while the branch network provides coefficients—akin to spectral expansions in Fourier or Laplace methods. Since spectral methods have already demonstrated strong performance in solving the Navier-Stokes equations, enforcing this structured decomposition allows Type 1 to better capture the underlying physics of flow evolution, outperforming the less structured feature fusion in Type 2.
To better understand the qualitative differences between these merge strategies, we visualize the predicted vorticity profiles along a vertical line at $x=0.5m$ for different time steps in Figure 7. These line plots show how vorticity values vary along the y-direction from 00 to $1m$, providing insight into each model’s ability to capture flow field features at different spatial locations. The visualization clearly demonstrates the limitations of the dot-product operation and the advantages of incorporating merge networks. At both time steps (Figure 7(a) and Figure 7(b)), the dot-product approach (without bias, which outperformed the with-bias variant) exhibits significant difficulties capturing local flow features, producing overly smoothed predictions that miss important vorticity fluctuations across the domain. In contrast, both Type 1 and Type 2 merge networks successfully reproduce the complex flow patterns, following the ground truth’s local peaks and valleys with notably higher fidelity.
(a) $t=0.5s$
(b) $t=0.75s$
Figure 7: Comparison of vorticity predictions along $x=0.5m$ between different merge strategies. The plots show vorticity values versus y-coordinate for two different time steps, comparing predictions from dot-product operation (without bias) and merge network approaches (Type 1 and 2) against ground truth (GT). Both merge network types demonstrate superior ability in capturing local flow features and fluctuations compared to the basic dot-product operation.
4.2 Effects of Spatial Point Sampling and Application of Automatic Mixed Precision
Building upon the optimal merge network configuration (Type 1) identified in Section 4.1, we investigate the impact of spatial point sampling and automatic mixed precision (AMP) training on model performance. Spatial point sampling refers to our approach of selecting only a subset of available spatial locations from the full 128×128 grid for training, leveraging DeepONet’s inherent capability to operate on arbitrary spatial points without requiring fixed grid structures. We explore three sub-sampling strategies: 16×16, 32×32, and 64×64 randomly selected points from the original 128×128 grid. AMP training refers to the strategic use of different numerical precisions (e.g., float32, float16) during model training to reduce memory usage and computational overhead without compromising model accuracy—a particularly relevant approach for DeepONet due to its intensive computational requirements in processing high-dimensional flow field data (for the details of AMP, please refer to Appendix A). For each sampling configuration, we conduct experiments both with and without AMP training to systematically evaluate their combined effects on computational efficiency and prediction accuracy. All other training settings remain consistent with Section 4.1.
Table 2: Comparison of different spatial sampling strategies with and without AMP training: Type 1 merge network is applied.
AMP Number of Sampled Points Training Hour Test Error
O 16×16 0.350 17.419
32×32 0.785 18.184
64×64 2.211 17.438
128×128 8.672 17.079
X 16×16 0.377 16.876
32×32 1.038 16.989
64×64 3.458 17.862
128×128 13.648 18.001
The results reveal important insights about DeepONet’s training efficiency through the combination of sampling strategies and AMP training. First, examining the impact of spatial point sampling, we observe that smaller sampling sizes can maintain prediction accuracy while significantly reducing computational costs. With AMP enabled, the 16×16 sampling strategy achieves comparable accuracy (test error: 17.419) to the baseline that uses all available points (test error: 17.079), while dramatically reducing the training time from 8.672 to 0.350 hours—a 96% reduction. Similar trends are observed in cases without AMP, where the 16×16 sampling rather achieves a test error of 16.876 compared to 18.001 for full sampling due to its effective sampling nature compared to naive all-point-sampling approach.
The comparison between AMP and non-AMP training reveals interesting trade-offs in computational efficiency and prediction accuracy across different sampling sizes. While non-AMP training with 16×16 sampling achieves slightly better accuracy (16.876 versus 17.419 with AMP) and requires only marginally longer training time (0.377 versus 0.350 hours), we opt for the AMP configuration with 16×16 sampling in subsequent experiments. This choice is motivated by two practical considerations: (1) our comprehensive study involves extensive experiments, each repeated multiple times to account for training stochasticity—in this context, even small improvements in computational efficiency become significant when scaled across numerous training runs; and (2) AMP’s computational benefits scale dramatically with sampling density—for the full 128×128 grid, AMP reduces training time by 36.5% compared to non-AMP training. This scalable efficiency gain underscores the importance of investigating AMP in DeepONet applications, as many practical implementations may require substantially denser spatial sampling than explored here. For these reasons—both to enhance efficiency across our numerous experimental runs and to leverage the scalable computational benefits that become increasingly significant with larger sampling sizes—we employ AMP in all subsequent experiments.
In Figure 8, we can observe the effects of different sampling resolutions on flow field prediction quality given AMP. The comparison demonstrates that even with the coarsest sampling resolution (16×16), the DeepONet successfully captures the essential vorticity patterns and intensity distributions present in the flow field at $t=0.25s$. As expected, the visual differences between sampling resolutions are minimal, with all configurations reproducing the key flow structures observed in the ground truth. This visual evidence supports our quantitative findings in Table 2, which indicate that the 16×16 sampling strategy achieves comparable prediction accuracy (MSE of 17.419) to the full 128×128 sampling (MSE of 17.079), while dramatically reducing the training time from 8.672 to 0.350 hours. The consistent visual quality across different sampling resolutions highlights DeepONet’s point-based prediction flexibility, which allows for efficient data utilization without substantial degradation in prediction performance.
Figure 8: To visually investigate the effects of number of sampled points, 4 cases with AMP in Table 2 are visualized: snapshots at $t=0.25s$. The comparison demonstrates that even with significantly reduced sampling (16×16), the model captures essential flow patterns similar to those observed in higher resolution samplings (32×32, 64×64) and the full dataset (128×128).
Note: The findings in this section highlight a fundamental advantage of DeepONet over grid-based architectures like CNNs or GNNs: its point-based prediction capability enables remarkable flexibility in data sampling, free from the constraints of fixed grid structures that limit traditional architectures and often require storing complete grid information in memory. This sampling flexibility, when combined with computational optimization techniques like AMP training, provides an effective strategy for enhancing DeepONet’s training efficiency while significantly reducing memory requirements compared to grid-based methods. Such adaptability is particularly valuable for real-world applications where computational resources may be limited, allowing users to strategically balance accuracy, memory usage, and computational cost through appropriate sampling strategies.
4.3 Effects of Positional Encoding
Building upon the optimal merge network configuration and sampling strategy identified in previous sections, merge Type 1 with 16×16 sampling with AMP, we investigate the impact of positional encoding (PE) on DeepONet’s prediction accuracy. PE transforms input coordinates using sinusoidal functions, potentially mitigating the spectral bias inherent in neural networks (Section 2.1.2). We systematically explore two distinct PE strategies: (1) PE($\mathbf{t}$): applying PE only to the temporal coordinate ($t$), and (2) PE($\mathbf{x,y,t}$): applying PE to all trunk network inputs ($x$, $y$, and $t$). For each strategy, we evaluate combinations of two key hyperparameters: the scale factor $\sigma$ controlling the frequency range, and the mapping size $m$ determining the dimension of encoded features. All other training settings remain consistent with Section 4.1.
Table 3: Effects of positional encoding strategies and hyperparameters on DeepONet performance
PE Strategy Scale Factor ($\sigma$) Mapping Size ($m$) Training Hour Test Error
Without PE - - 0.350 17.419
PE($\mathbf{t}$) 1 16 0.355 18.271
32 0.363 16.751
64 0.363 17.331
5 16 0.355 17.797
32 0.365 17.578
64 0.359 18.065
10 16 0.358 16.619
32 0.368 16.100
64 0.357 17.421
PE($\mathbf{x,y,t}$) 1 16 0.356 21.709
32 0.355 24.981
64 0.352 24.197
5 16 0.354 43.161
32 0.359 32.261
64 0.351 29.300
10 16 0.358 81.884
32 0.356 38.925
64 0.306 NaN*
    Four out of five models diverged, indicating significant instability in this setting due to high-frequency embedding at spatial coordinates.
The comparison of different PE strategies reveals several important insights about DeepONet’s behavior with coordinate encoding (Table 3). First, it’s noteworthy that all models demonstrate similar training times (approximately 0.35-0.37 hours with NVIDIA A6000 GPU), indicating that the application of PE does not meaningfully impact computational efficiency. This consistent training overhead across all configurations makes PE an attractive option for potential performance improvement without computational penalties. The results show a clear distinction between temporal-only and full spatio-temporal encoding strategies. PE($\mathbf{t}$) improves model performance, with the best configuration ($\sigma=10$, $m=32$) achieving a test error of 16.100 compared to 17.419 without PE—a 7.57% improvement. In contrast, PE($\mathbf{x,y,t}$) leads to deteriorated performance across all configurations, with test errors increasing significantly as the scale factor increases. This deterioration becomes particularly severe with higher scale factors, culminating in training instability at $\sigma=10$ and $m=64$ where four out of five training runs diverged, indicating that high-frequency embeddings of spatial coordinates can destabilize the training process in our case.
The impact of hyperparameters reveals important considerations for PE implementation. The scale factor $\sigma$, which controls the frequency range of the encoding, shows strong influence on model performance. In PE($\mathbf{t}$), higher $\sigma$ values generally lead to better performance, suggesting that temporal dynamics in our flow field benefit from higher-frequency encodings. However, this relationship between performance and $\sigma$ is highly dependent on the underlying temporal characteristics of the flow field, emphasizing the importance of careful hyperparameter tuning for specific applications. Furthermore, the mapping size $m$ demonstrates a non-monotonic relationship with model performance. For instance, in PE($\mathbf{t}$), increasing $m$ from 32 to 64 consistently degrades performance across all scale factors, indicating that larger mapping sizes do not necessarily translate to better predictions. This suggests an optimal intermediate dimension for the encoded features that balances expressiveness with model complexity, with $m=32$ proving to be the optimal value in our case.
In conclusion, our results demonstrate that positional encoding, when properly configured and applied selectively to temporal coordinates, can enhance DeepONet’s prediction accuracy without compromising computational efficiency. The 7.57% improvement achieved with PE($\mathbf{t}$) requires no additional architectural complexity or significant implementation effort, making it a valuable addition to DeepONet implementations for spatio-temporal prediction tasks. However, practitioners should carefully consider both the encoding strategy and hyperparameter selection, since inappropriate choices—particularly in spatial coordinate encoding—can significantly impair model performance.
5 Application of Multi-Fidelity DeepONet
5.1 Effects of Different Transfer Learning Approaches with Different Low-Fidelity Dataset
Building upon our optimized DeepONet architecture with temporal positional encoding ($\sigma=10$, $m=32$) in Section 4.3, we evaluate the following various MF-DeepONet architectures, all designed with the same model parameter counts to ensure fair comparison:
  1. 1.
    Lu et al. [18] (conventional MF-DeepONet): while the original MF-DeepONet approach proposed by Lu et al. [18] combines both residual learning and input augmentation within a coupled two-DeepONet framework, we strategically implement only the input augmentation aspect. We discard the coupled use of two DeepONet models and residual learning component due to their significant limitation: they fundamentally require identical query points across fidelity levels, which severely restricts the flexible usage of multi-fidelity datasets. Therefore, its implementation includes the sequential architecture where LF and HF DeepONets are trained separately as our proposed MF-DeepONet framework, with the LF model’s output directly feeding into the HF model’s trunk net—input augmentation aspect proposed by Lu et al. [18]. This architecture still necessitates that both networks execute in sequence during all training and inference operations, requiring much more computational time than our proposed framework.
  2. 2.
    Proposed: as proposed in Section 2.2.1, our MF-DeepONet employs fine-tuning where the pre-trained LF DeepONet’s branch and trunk networks and frozen, while only the merge network parameters are updated during HF training. This approach preserves valuable LF representations while adapting specifically to HF features.
  3. 3.
    Full-tuning: all network components (branch, trunk, and merge networks) are simultaneously trainable, potentially compromising the established low-fidelity knowledge while predominantly focusing on high-fidelity feature acquisition.
  4. 4.
    Linear probing: takes a highly restrictive approach by freezing all layers except the final layer of the merge network.
We systematically evaluate these approaches across different LF data configurations, varying both spatial resolution (16×16, 32×32, 64×64) and dataset size (50, 100, 200, 300 samples)—note that HF dataset size is fixed as 100. For fair comparison with the single-fidelity DeepONet which uses 800 training epochs, both LF and HF training phases in the multi-fidelity approaches are set to 400 epochs each, ensuring the total number of training epochs remains constant across all experiments in this paper.
The results, presented in Table 4 and also visualized as Figure 9 for easier comparison of MSE and training hour, reveal several significant insights about MF-DeepONet performance. First, our proposed fine-tuning approach consistently outperforms all other methods, including the input augmentation implementation of MF-DeepONet from Lu et al. [18]. The conventional MF-DeepONet approach (Lu et al. [18]) shows limited effectiveness, with MSE values ranging from 25.407 to 26.388 for 16×16 resolution data, significantly underperforming even single-fidelity training (MSE 16.100). While these inferior results may partly stem from our necessary omission of the residual learning component from the original Lu et al. [18] framework, this design choice was deliberate and essential for our study’s objectives. The residual learning approach, while potentially beneficial in certain contexts, fundamentally restricts DeepONet’s inherent data flexibility by requiring identical sampling points across fidelity levels—a constraint that severely limits practical applicability in real engineering scenarios where sampling locations often differ between low and high-fidelity simulations. Even at higher resolutions (64×64 LF data with 300 samples), this approach achieves an MSE of only 15.940, which is 76% higher than our fine-tuning method’s 9.057 MSE. Additionally, the sequential architecture of Lu et al. [18]’s approach—requiring both LF and HF DeepONet networks to be executed in sequence during inference—leads to consistently higher computational costs, with training times approximately 25-35% longer than our proposed method across all configurations.
(a) MSE
(b) Training hour
Figure 9: Performance comparison of conventional DeepONet (Lu et al. [18]), fine-tuning, full-tuning, and linear probing approaches for MF-DeepONet. Results show MSE and training time across different LF data sizes (50-300) and spatial resolutions (16×16: dotted, 32×32: dashed, 64×64: solid lines). This figure illustrates a visual representation of the numerical results presented in Table 4, enabling direct comparison of performance metrics across different model configurations.
Table 4: Comparison of different MF-DeepONet architectures. The values highlighted in gray serve as baseline performance metrics for the analysis of time-derivative guided sampling efficiency in Section 6.
Architecture LF Data Resolution LF Data Size Training Hour MSE
Single-fidelity 0.368 16.100
Lu et al. [18] (conventional MF-DeepONet) 16×16 50 0.348 26.273
100 0.411 26.388
200 0.492 26.123
300 0.598 25.407
32×32 50 0.341 27.029
100 0.399 23.432
200 0.513 20.694
300 0.630 18.936
64×64 50 0.390 25.710
100 0.407 21.053
200 0.518 16.628
300 0.633 15.940
Proposed 16×16 50 0.219 19.494
100 0.268 17.739
200 0.370 15.238
300 0.466 15.096
32×32 50 0.213 19.336
100 0.269 14.324
200 0.362 11.791
300 0.460 10.536
64×64 50 0.221 17.673
100 0.272 14.698
200 0.369 10.782
300 0.469 9.057
Full-tuning 16×16 50 0.251 21.687
100 0.301 21.119
200 0.401 17.893
300 0.504 17.974
32×32 50 0.254 22.050
100 0.303 16.807
200 0.416 14.772
300 0.506 13.146
64×64 50 0.256 20.870
100 0.316 17.208
200 0.414 13.612
300 0.538 11.171
Linear probing 16×16 50 0.176 73.678
100 0.225 64.520
200 0.332 60.894
300 0.433 61.179
32×32 50 0.175 48.306
100 0.221 36.800
200 0.324 32.190
300 0.429 29.881
64×64 50 0.181 35.587
100 0.239 26.301
200 0.350 21.373
300 0.466 20.163
Second, fine-tuning consistently outperforms both full-tuning and linear probing across all configurations, demonstrating the effectiveness of our strategy to preserve learned low-fidelity latent representations by branch/trunk networks while adapting only the merge network to high-fidelity features. Particularly noteworthy is that even with the coarsest LF resolution (16×16), fine-tuning achieves better performance compared to single-fidelity training (MSE of 15.238 vs 16.100) when provided with sufficient LF data (200 samples). This improvement is remarkable given that the 16×16 resolution data exhibits significantly different flow physics compared to higher resolutions, as evidenced in Figure 6. While full-tuning provides more flexibility through complete network adaptation, this additional freedom may not be optimal, as it can lead to less efficient use of valuable low-fidelity feature representations learned during pre-training. The superior performance of fine-tuning suggests that the branch and trunk networks effectively capture essential parameter and spatio-temporal representations during low-fidelity training, which can be successfully leveraged for high-fidelity predictions through careful adaptation of the merge network alone.
In terms of LF dataset, as expected, increasing the resolution of LF data from 16×16 to 64×64 leads to consistent improvements in prediction accuracy across all transfer learning approaches. With 300 LF samples at 64×64 resolution, fine-tuning achieves an MSE of 9.057, representing a 43.7% improvement over single-fidelity training (where MSE is 16.001). Figure 10 presents a comparative visualization of vorticity field predictions at two different time steps ($t=0.5s$ and $t=0.75s$). Our proposed MF-DeepONet with fine-tuning demonstrates notably superior prediction accuracy compared to the single-fidelity approach, particularly in regions highlighted by white dotted boxes. Within these regions, the single-fidelity model exhibits significant deviations from the ground truth, failing to capture the correct vorticity patterns and flow structures. In contrast, our proposed MF-DeepONet successfully reproduces the complex flow features, maintaining fidelity to the true flow dynamics across both temporal snapshots.
(a) $t=0.5s$
(b) $t=0.75s$
Figure 10: Visual comparison of the predicted vorticity field at $t=0.5s$ and $t=0.75s$ between single-fidelity DeepONet and multi-fidelity DeepONet with fine-tuning (with LF data resolution of 64×64 and data size of 300) within test scenario.
Note: The results in Table 4 reveal important insights about knowledge transfer between fidelity levels. Linear probing consistently performs poorly across all configurations, with MSE values substantially higher than both fine-tuning and single-fidelity approaches. This underperformance is scientifically significant—it confirms the substantial differences between low-fidelity and high-fidelity flow physics, as the frozen representations from low-fidelity models inadequately capture high-fidelity features without adaptation. Despite these fundamental differences, our fine-tuning approach demonstrates remarkable effectiveness—even when using coarse 16×16 resolution low-fidelity data with markedly different physical characteristics from ground truth physics, the model still outperforms training on HF data alone. Specifically, fine-tuning with 16×16 resolution low-fidelity data achieves MSE values of 15.238 (with 200 samples) and 15.096 (with 300 samples), representing improvements of 5.4% and 6.2% respectively over the single-fidelity baseline (MSE: 16.100). This confirms that our selective adaptation approach effectively leverages useful information from low-fidelity simulations, even when they exhibit partially different physical characteristics compared to their high-fidelity counterparts.
5.2 Performance Analysis with Limited High-Fidelity Data
To further investigate the potential of our MF-DeepONet framework in scenarios where HF data is scarce, we conduct a systematic study using reduced HF dataset sizes. While our previous analysis in Section 5.1 utilized 100 HF samples, here we examine the model’s performance across a range of smaller HF dataset sizes: 20, 40, 60, and 80 samples. This investigation addresses a critical challenge in real-world engineering applications, where HF data is typically extremely limited. Based on our findings from Section 5.1, we focus on the more effective LF resolutions of 32×32 and 64×64, excluding the 16×16 resolution due to its relatively limited performance benefits. For each LF resolution, we explore two different LF dataset sizes (200 and 300 samples) to understand how the quantity of LF data affects the model’s ability to compensate for limited HF data. This systematic approach will allow one to establish practical guidelines for efficiently allocating computational resources between LF and HF simulations in practical applications.
The results shown in Figure 11 reveal several important insights about the MF-DeepONet’s performance with limited HF data. First, examining the MSE trends (Figure 11(a)), we observe a consistent pattern of improved prediction accuracy as the number of HF samples increases across all configurations. The 64×64 resolution configurations consistently outperform their 32×32 counterparts, with the 64×64 with 300 samples combination achieving the best performance across all HF dataset sizes. A particularly noteworthy observation is that increasing the LF dataset size from 200 to 300 samples yields more improvements in accuracy for the 64×64 resolution compared to the 32×32 resolution. For instance, with 60 HF samples, the 64×64 resolution shows a reduction in MSE from 13.476 (200 LF samples) to 11.364 (300 LF samples)—a 15.7% improvement—while the 32×32 resolution exhibits a smaller improvement from 16.031 to 13.760, representing a 14.2% reduction in error.
(a) MSE
(b) Training hour
Figure 11: MF-DeepONet performance with limited high-fidelity data: different resolutions of LF data with different LF data sizes are explored. The label “LF 32×32 (200)” represents a case where LF data is from a 32×32 resolution with a dataset size of 200—all other labels follow the same notation, indicating the corresponding LF resolution and dataset size.
The training time analysis (Figure 11(b)) reveals an expectable trade-off. Configurations with 300 LF samples consistently require longer training times compared to their 200 LF sample counterparts due to the longer training time required in LF deepONet. However, this increased computational cost appears justified by the improved accuracy, particularly for the 64×64 resolution cases.
6 Time-Derivative Guided Sampling for Enhancing Multi-Fidelity DeepONet
6.1 Effects of Time-Derivative Guided Sampling
Based on our analysis in Section 5.2, we now investigate two optimal configurations of our fine-tuned MF-DeepONet framework (highlighted in Figure 11(a) with yellow circles and black dashed edges) that achieve an exceptional balance between prediction accuracy and high-fidelity data efficiency:
  1. 1.
    Case 1: MF-DeepONet with 32×32 LF resolution (300 samples) trained using 80 HF samples
  2. 2.
    Case 2: MF-DeepONet with 64×64 LF resolution (300 samples) trained using 60 HF samples
For these two configurations, we analyze the effectiveness of our time-derivative guided sampling strategy in practice. This physics-guided approach, formalized in Algorithm 1 in Section 2.2.2, strategically identifies regions with pronounced temporal dynamics from the pre-trained LF DeepONet, enabling more efficient HF data collection in physically meaningful areas.
In our experiments, we set the sampling ratio $r=0.1$, meaning that 10% of the HF spatial points are selected based on the temporal derivative information from LF DeepONet, while the remaining 90% are sampled uniformly at random to maintain spatial coverage. When compared to their respective baselines only with uniform sampling (without time-derivative guided sampling), both configurations show substantial improvements in prediction accuracy, as detailed in Table 5. The 32×32 LF resolution case shows a 10.02% reduction in MSE (from 11.759 to 10.581), while the 64×64 LF resolution case achieves a significant MSE reduction of 20.73% (from 11.364 to 9.008). These accuracy gains come with moderate increases in computational cost—7.68% and 10.74% increases in training time for Case 1 and Case 2, respectively. This additional computational overhead, primarily stemming from the automatic differentiation calculations required for temporal derivatives, is well justified given the significant improvement in accuracy. The greater improvement in the 64×64 configuration indicates that higher-resolution LF data produces a more accurate LF DeepONet model, which in turn provides more reliable temporal derivative information. This enhanced precision in identifying dynamically important regions results in more effective HF data sampling, improving the utilization of limited high-fidelity resources.
Table 5: Performance comparison between uniform (Before) and time-derivative guided sampling strategies (After).
Metric Case Before After Change
MSE Case 1: LF 32×32 w/ 80 HF 11.759 10.581 -10.02%
Case 2: LF 64×64 w/ 60 HF 11.364 9.008 -20.73%
Training hour Case 1: LF 32×32 w/ 80 HF 0.469 0.505 +7.68%
Case 2: LF 64×64 w/ 60 HF 0.475 0.526 +10.74%
The above quantitative improvements are further supported by the qualitative analysis shown in Figure 12, which compares vorticity profiles at different time steps. At $t=0.25s$ (Figure 12(a)), all models show similar prediction capabilities, closely following the ground truth profile. This comparable performance in the early stage is expected, as the flow field remains relatively simple and closer to the initial conditions. However, at $t=0.5s$ (Figure 12(b)), clear differences emerge in the models’ ability to capture complex flow dynamics. This later time step exhibits significantly more complicated flow development with pronounced temporal variations—precisely the conditions where our time-derivative sampling strategy should excel. Indeed, particularly in the middle region ($y=0.3m$ to $0.7m$), models trained with time-derivative sampling (presented as dashed lines) more accurately reproduce the local fluctuations and subtle variations in the vorticity profile. This superior performance demonstrates that our approach automatically identifies and focuses computational resources on regions with significant temporal dynamics, enhancing prediction accuracy exactly where conventional methods struggle. In contrast, models without sampling tend to oversimplify these local features, capturing only the overall trend while missing important flow structure details. These results confirm that our physics-guided sampling strategy effectively targets time-sensitive regions of the flow field ($t=0.5s$ than $t=0.25s$), showing particular advantage in predicting complex time-dependent dynamics that develop as the simulation progresses.
(a) $t=0.25s$
(b) $t=0.5s$
Figure 12: Comparison of vorticity profiles from ground truth (GT) and MF-DeepONet predictions with and without time-derivative guided sampling. At early stages ($t=0.25s$), all models show comparable performance, but at later times ($t=0.5s$), models with time-derivative sampling demonstrate superior accuracy in capturing complex flow dynamics. Case 1: 32×32 LF (300 samples) with 80 HF samples; Case 2: 64×64 LF (300 samples) with 60 HF samples.
6.2 Reduced High-Fidelity Data Requirement through Time-Derivative Guided Sampling
The time-derivative guided sampling strategy demonstrates remarkable effectiveness, achieving superior predictive accuracy compared to conventional uniform sampling when using the same amount of HF data (Section 6.1). Therefore, in this section, we investigate how this approach can reduce the required HF dataset size while maintaining comparable accuracy to models trained without time-derivative sampling—quantifying the potential data efficiency gains enabled by our physics-guided sampling method. For MF-DeepONet using 64×64 LF resolution data (Case 2), our sampling approach achieves an MSE of 9.008 with only 60 HF samples, surpassing the performance of models trained on the full 100 HF samples with uniform sampling (MSE 9.057, marked as gray in Table 4). This achievement is particularly significant as it represents both improved accuracy and a 40% reduction in required HF data. Similarly, with 32×32 LF resolution data (Case 1), our time-based sampling achieves prediction accuracy (MSE 10.581) using only 80 HF samples, effectively matching the performance of training with the complete HF dataset and uniform sampling (MSE 10.536, marked as gray in Table 4).
Figure 13 provides a visual comparison between the uniform-sampling-based multi-fidelity model trained with the 100 HF dataset and the multi-fidelity model coupled with time-derivative guided sampling approach using only 60 HF dataset. At both time steps ($t=0.5s$ and $t=0.75s$), we observe that both models produce visually comparable flow field predictions that closely match the ground truth. The error distributions demonstrate that, despite using significantly less HF data (only 60%), the time-derivative sampling approach maintains similar prediction quality to the uniform-sampling model. While there are subtle differences in the error patterns—particularly visible in certain regions at $t=0.75s$ where the proposed physics-guided sampling approach shows slightly higher localized errors—these minor differences in error patterns are insignificant when considering the substantial 40% reduction in HF data requirements achieved by our sampling approach. This visual evidence reinforces our quantitative findings that strategic selection of high-fidelity training points based on temporal dynamics enables efficient utilization of limited HF data with minimal degradation in prediction quality.
(a) $t=0.5s$
(b) $t=0.75s$
Figure 13: Visual comparison of the predicted vorticity field at $t=0.5s$ and $t=0.75s$ between two models: 1) baseline uniform-sampling-based MF-DeepONet with LF 64×64 and 100 HF samples, and 2) MF-DeepONet with LF 64×64 and only 60 HF samples guided by time-derivative sampling. Despite using 40% fewer HF samples, the model with time-derivative sampling maintains comparable prediction quality, with only minor differences in error patterns.
Figure 14 presents an analysis of error behavior of the test data over time for Cases 1 and 2, comparing scenarios with and without time-derivative guided sampling. It is clear from the plot that incorporating time-derivative guided sampling significantly mitigates error increase as temporal prediction progresses for both cases. This highlights that the physics-guided, time-derivative sampling not only enhances accuracy at individual time steps but also substantially limits error propagation over extended prediction intervals. Such improved long-term predictive stability is particularly valuable for industrial applications where accuracy over prolonged simulation times is crucial.
Figure 14: Comparison of MSE behavior over time steps (200 steps stands for a second) for Case 1 and Case 2, with and without time-derivative guided sampling. The results clearly demonstrate that the time-derivative guided sampling effectively reduces error over time, thereby significantly improving prediction stability and accuracy.
Note: While our sampling strategy leverages temporal derivatives as the key physical indicator, the MF-DeepONet framework itself is not confined to this specific sampling approach. Temporal derivatives serve as an effective physics-informed criterion for identifying critical regions in the examples studied here, capturing fundamental characteristics where rapid changes often indicate important physical transitions. However, the modular nature of our framework allows for integration with various alternative sampling techniques tailored to specific applications. For instance, future work could explore uncertainty-driven sampling for stochastic systems. This flexibility represents a significant advantage of our approach—the underlying MF-DeepONet architecture can be coupled with domain-specific sampling strategies while maintaining its core multi-fidelity benefits. This opens numerous avenues for researchers to develop specialized sampling techniques optimized for their particular physical systems, potentially yielding even greater efficiency gains beyond what we’ve demonstrated with temporal derivatives in this study.
7 Conclusion
This study presents a novel multi-fidelity DeepONet (MF-DeepONet) framework for efficient spatio-temporal flow field prediction with reduced high-fidelity data requirements. Through strategic architectural enhancements and physics-guided sampling, we address critical challenges in operator learning for fluid dynamics applications. Our merge network architecture enables more complex feature interactions between operator and coordinate spaces, achieving a 50.4% reduction in prediction error compared to traditional dot-product approaches. The implementation of temporal positional encoding and efficient point-based sampling strategies further improves prediction accuracy by 7.57% while dramatically reducing training time by 96%. The transfer learning-based multi-fidelity framework leverages knowledge from pre-trained low-fidelity models to guide high-fidelity predictions, demonstrating a 43.7% improvement in accuracy compared to single-fidelity training. Most significantly, our time-derivative guided sampling strategy demonstrates two key advantages: (1) it reduces prediction error by 20.7% compared to conventional uniform sampling when using the same amount of high-fidelity data, and (2) it requires only 60% of the high-fidelity data to achieve comparable accuracy to models trained with full data using uniform sampling—a crucial advancement for computationally expensive simulations.
Despite these promising results, our framework has several limitations that warrant further investigation. Although highly effective for our flow problems, the time-derivative guided sampling strategy requires broader validation across diverse physical systems with different temporal characteristics. Additionally, our current transfer learning approach—which completely freezes branch and trunk networks—may be overly restrictive; more flexible layer-freezing strategies based on data similarity metrics could potentially yield further improvements. Furthermore, the computational overhead introduced by automatic differentiation in temporal derivative calculations presents opportunities for developing more efficient sampling implementations. Lastly, to enhance practical applicability in industrial settings, extending the framework to accommodate irregular domains, complex geometries, and moving boundaries represents an important direction for future work.
Future research should focus on developing adaptive physics-guided sampling criteria beyond temporal derivatives to improve cross-domain generalizability. An automated layer-wise similarity analysis between low-fidelity and high-fidelity features could enable more nuanced transfer learning approaches with selective network component freezing. Computational efficiency could be enhanced through surrogate-based derivative estimation or alternative physics-guided sampling metrics that avoid expensive automatic differentiation calculations. Finally, incorporating geometric information directly into the sampling strategy would enable effective point selection in complex domains, broadening the framework’s applicability across diverse scientific disciplines where high-fidelity data acquisition remains challenging. These advancements would further establish MF-DeepONet as a practical solution for real-time flow prediction in data-constrained environments.
Acknowledgments
This work was supported by the National Research Foundation of Korea (2018R1A5A7025409), and the Ministry of Science and ICT of Korea grant (No. 2022-0-00969, No. 2022-0-00986, and No. RS-2024-00355857), the Ministry of Trade, Industry & Energy grant (RS-2024-00410810), and the National Research Council of Science & Technology (NST) grant by the Korea government (MSIT) (No. GTL24031-000)
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Appendix A Automatic Mixed Precision (AMP)
AMP is a method that blends half-precision and full-precision computations to accelerate training and reduce memory usage, while maintaining stability. In simpler terms, it uses smaller data types (for example, 16-bit floats) during most calculations for speed and efficiency, then switches to larger data types (like 32-bit floats) for operations that risk losing too much detail. This strategy lowers the amount of memory required and speeds up math operations on modern hardware.
A crucial part of AMP is a safeguard mechanism that dynamically scales gradients to avoid floating-point overflow or underflow. In practice, a scaling factor is applied to gradients during backpropagation. After the updates are computed, this factor is reversed before adjusting the model’s parameters. If the training process detects an overflow—meaning the gradients became too large—the scaling factor is automatically reduced for the next round of updates, preserving numerical stability. Additional insights into mixed precision and gradient scaling can be found through sources such as the NVIDIA Developer Blog [24].